# Identifying Halos in Cosmological Simulations with Continuous Wavelet Analysis: The 2D Case

### v2

@@ 2.3 SIMBA Simulation and Pseudo-2D Data
SIMBA (Davé et al., 2019) is a meshless finite mass hydrodynamic cosmological simulation based on the GIZMO cosmological gravity plus hydrodynamics solver (Hopkins, 2015), assuming $h=0.68$, $\Omega_{\rm\Lambda}=0.7$, $\Omega_{\rm m}=\Omega_{\rm dm}+\Omega_{\rm b}=0.3$, $\Omega_{\rm b}=0.048$, $n_{s}=0.97$ and $\sigma_{8}=0.82$, which are the Planck cosmological parameters (Planck Collaboration et al., 2016). This simulation consists of several sub-simulations with different box lengths, particle numbers and even different physics, e.g. no AGN, no jet and pure dark matter.
The purpose of this paper is to identify halos from particle data, without delving into the detailed physics of the simulation. We therefore use the m50n512 dark matter sub-simulation, which is a pure dark matter run without any complicated processes such as star formation, gas cooling, feedback and so on. This sub-simulation has a box length of $50h^{-1}\rm Mpc$ and contains $512^{3}$ dark matter particles.
To obtain 2D data, we select the snapshot of redshift $z=0$ from this sub-simulation and cut a slice in the z-direction, and keep only the x-, y- coordinates of these particles. This is actually pseudo-2D data with the dimension reduced from the original three-dimensional (3D) data, but it still contains a lot of structures and is sufficient to test our identification algorithm.
We use four slices of pseudo-2D data at ${\rm z}=11.4$, $28.4$, $38.0$ and $45.7\,{h^{-1}{\rm Mpc}}$ with a thickness of $\Delta{\rm z}=0.1\,{h^{-1}{\rm Mpc}}$, containing 323,716, 228,187, 303,692 and 312,372 particles, respectively. The CIC grid of these slices are shown in Figure 1.
@@ 3.1 Procedure of Identification
(1) Calculation of the grid CWT at different scales using a grid-assigning method, as shown in Equation (6), results in an $\mathcal{O}(N)$ time complexity.
(2) Determination of thresholds to filter those local maxima generated by Poisson noise at different scales, as discussed in Section 3.3, is performed using either the Monte Carlo method or a pre-calculated empirical formula.
@@
In the following, we present the detailed calculations for each step of the process.
Figure 3: Left: Threshold-$k_{w}$ relationship at different particle numbers, ranging from $N=10^{5}$ (bottom, black) to $10^{7}$ (top, blue). We can observe that the relationship between the variables demonstrates high linearity. Right: The slope and intercept of the linear relation from the left panel are shown as functions of the particle number.
@@ 3.2 Grid-based CWT Computation
In Equation (6), we have derived that the grid CWT is the sum of the wavelets centered on the position of the particles. Since the MH wavelet $\Psi(x)$ in Equation (8) has an approximate compact support, it affects only a finite range near its center. For some $r_{\rm c}>0$, if $|\Psi(|\mathbf{x}|>r_{\rm c})|$ is less than, say, $10^{-10}$, then we can truncate $\Psi(|\mathbf{x}|>r_{\rm c})$ to be zero. This property allows a convenient calculation of the grid CWT, since a wavelet of a given scale will only have a non-zero value at grid points near its center.
@@
In this way, $k_{w}=1$ always represents an intermediate scale when identifying halos, and for the current work, $k_{w}=1$ corresponds to a scale factor of $20\,{h{\rm Mpc}^{-1}}$. A dynamic mesh is created based on the variable $k_{w}$ to save memory and speed up computation. To accurately compute these translated wavelets on the grid, the grid point closest to the position of the particle is selected as the center of the grid wavelet, and this grid point is extended by ${\rm Int}(10/k_{w})+10$ grid points in each direction. We will discuss the resolution parameters in the above two equations in detail in the next section.
The $N_{\rm g}^{2}$ grid is initialized with zeros, and each individual CWT, corresponding to its particle in the $N_{\rm W}^{2}$ grid, is calculated and added to the appropriate part of the $N_{\rm g}^{2}$ grid. This part is centered at the grid point nearest to the particle and extends ${\rm Int}(10/k_{w})+10$ grid points in each direction. If a particle is located near the edge of the simulation box, its extended ${\rm Int}(10/k_{w})+10$ grid points may overflow beyond the $N_{\rm g}^{2}$ grid. In this scenario, the overflowed grid points is moved to the opposite side, using the periodic boundary condition applied to the particle data.
Figure 4: The largest halo within the slice at ${\rm z}=38.0\,{h^{-1}{\rm Mpc}}$, presented from left to right at low ($250500\times k_{w}+$ 250), medium ($5001000\times k_{w}+$ 500), and high ($10002000\times k_{w}+$ 1000) resolutions.
@@ 3.3 Cross-Scale Analysis of Local Maxima and Thresholding
After calculating the CWT grid, the clustering of particles can be easily analyzed (Slezak et al., 1990). When using wavelets to identify haloes, it is generally agreed that each halo should be located at the position of its corresponding local maximum. The scale at which we find the maximum represents the size of the structure (e.g. Bendjoya et al., 1991; Flin & Krywult, 2006; Grebenev et al., 1995; Slezak et al., 1993; Cayón et al., 2000; Lazzati et al., 1999; Patrikeev et al., 2006; Hayn et al., 2012). Since our wavelet is not orthogonal, finding its local maximum directly in 3D space (2D location plus 1D scale parameter) can be misleading, as a single structure may reach its maximum at multiple scales (Hayn et al., 2012; Ellien et al., 2021). Therefore, we need to check these maxima across scales to ensure that we identify the maximum that corresponds to the actual structure, that is, the largest maxima near the position of this structure.
@@
Also, the particle distribution represents a Poisson sampling of the so-called real density field, which introduces Poisson noise that can lead to spurious local maxima when calculating the CWT using these particle data (Bijaoui et al., 1992). To mitigate this issue, a significance test should be conducted to assess the likelihood of detecting the true structure as opposed to Poisson noise. We apply thresholds obtained from Monte Carlo simulations and eliminate any maxima that do not meet these thresholds across each scale. A detailed description of this process will be presented in the subsequent section.
@@
Once we have identified all thresholded local maxima, we divide the simulation box into an additional set of the $N_{\rm cs}^{2}$ grid for cross-scale comparison. Our cross-scale comparison is based on a grid approach, not only because of the non-orthogonality of the wavelet but also due to the use of a dynamic mesh. The introduction of the new $N_{\rm cs}^{2}$ grid provides a consistent spatial resolution across scales, rather than varying resolutions.
@@
We then convert the location of these maxima, which is expressed as the count of grid lengths at their corresponding scale, to the position on the $N_{\rm cs}^{2}$ grid, starting from the larger scale to the smaller scale (from a smaller $w$ to a larger $w$). When we map a maximum onto the $N_{\rm cs}^{2}$ grid for cross-scale comparison, we compare its CWT value with the values of the surrounding $3\times 3$ grid points centered on the grid point of this maximum. If this maximum has a higher CWT value than all nine values within the surrounding grid points, we clear the information stored in those nine grid points and record the information of this maximum in its corresponding grid point. This ensures that there is only one maximum with the highest CWT value within every $3\times 3$ grid block of the $N_{\rm cs}^{2}$ grid. These maxima correspond to halo candidates in our method.
Figure 5: The 2D CWT of the slice of ${\rm z}=38\,{h^{-1}{\rm Mpc}}$ with four different scales, ranging from $2.29\,{h{\rm Mpc}^{-1}}$ to $35.24\,{h{\rm Mpc}^{-1}}$. The sizes of the structures presented in these CWTs vary depending on the scale parameter, extending from large to small scales.
@@ 3.4 Grid Segmentation and Halo Boundary Delineation
To identify a halo, a clear boundary must be established between the halo and the background universe. Fortunately, the wavelet method provides a natural boundary through its inherent properties. If the signal’s shape is consistent with that of the wavelet, such as at the peak of the signal, then a positive CWT value is obtained in that region. Conversely, if the signal shape is opposite to the wavelet, such as at the valley of the signal, then a negative CWT value is obtained in that region. Besides that, the matching between the peak and the negative ring of our wavelet also leads to negative CWT value (see Figure 5). No doubt, the positive CWT value inside the peak, while the negative CWT value, where the peak and the negative ring matches, outside the peak. If a wavelet matches with a peak in a positive signal, that means the positive part has the similar size as the peak, leaving the negative ring outside. In this way, points where the CWT equals zero is an intermediate position and serve as the boundary of the peak in the CWT field. However, since the CWT is calculated on a finite grid, finding the grid points where the CWT equals zero can be insufficient. In some cases, there may not be any grid points that are strictly equal to zero.
@@
Fortunately, we can also use all the outer regions as boundaries to isolate halos. This allows us to segment the full CWT grid at each scale into ‘islands’ of positive values surrounded by a ‘sea’ of zero or negative values. These ‘islands’ represent clusters at that scale. Similar segmentation methods have been employed in many previous studies to identify various structures within the CWT grid (e.g. Baluev & Rodionov, 2020; Freeman et al., 2002), as well as in the DWT grid (e.g. Pagliaro et al., 1999; Mertens & Lobanov, 2015).
@@
This method is effective for small and medium scales, but it is not suitable for large scales. The support of wavelet is more extensive at larger scales, which results in the linking of nearby structures by ‘positive bridges’. Additionally, the lower grid resolution at larger scales exacerbates this situation. Regardless of the constant boundary condition applied, some structures are invariably linked by a ‘positive bridge’.
@@
Despite its substructures, a gravity-bound halo exhibits a unimodal structure. At a larger scale, these substructures have a negligible impact on the CWT. Therefore, any actual halo must have a unimodal CWT. With this in mind, we introduce a new boundary condition akin to that described in (Freeman et al., 2002), which divides closely situated structures by identifying the saddle point between them. At any scale, we start by identifying all local maxima in the CWT and use these as the basis for initial segmentation. During each iteration, we examine all connected grid points to determine if any are lower than their connected structure grid points yet higher than zero. We then expand these structures to include neighboring grid points that meet the aforementioned criteria. The expansion of structures continues until no neighboring grid points can be incorporated. At that point, our segmentation process is complete.
@@
As mentioned in Section 3.3, a single structure can generate multiple ‘islands’ across various scales. Therefore, it is necessary to associate each halo candidate maximum with an ‘island’ to prevent multiple detections. Additionally, substructures must be excluded, as they can also correspond to local maxima and ‘islands’, which can lead to confusion when identifying halos. We zoom in on all the ‘islands’ and their corresponding background ‘sea’ at different scales to the highest resolution, which is the resolution of the smallest scale factor $w$. Every halo candidate is then assessed from larger to smaller scales. If a halo candidate is located within an ‘island’ at its scale and this ‘island’ does not overlap with any confirmed structures, it is registered as a halo.
Figure 6: The CWT halos of our pseudo-2D data. The slices are arranged in the order of ${\rm z}=11.4$, $28.4$, $38.0$ and $45.7\,{h^{-1}{\rm Mpc}}$ from top left to bottom right.
@@ 3.5 Refinement and Smoothing of Halo Boundaries
After registering all halos on a map with the highest resolution (smallest scale), we examine each particle to determine if it is located within a halo and label it with the corresponding halo index. However, due to our limited grid resolution, the CWT halo in the densest regions has very sharp edges. To address this issue, we propose checking if the particle is located within any particular halo grid, as shown in left panel of Figure 2. To prevent sharp edges, we examine the halo attribution of the four nearest grid points for each particle and calculate the particle’s weight to the four grid points using a CIC weighting formula, as
@@
where ${\rm x}_{i}$, ${\rm y}_{j}$ represent the normalized distances to the $i,j$-th grid in x- and y-direction, respectively, and $i,j\in[1,2]$ denote the four nearest grid points. After obtaining the weights, we sum them up if some grid points are assigned to the same halo. The particle is then assigned to the halo with the largest weight (we consider the background as halo ‘0’). From Figure 2, we can see that after weighting, the previously sharp edges are now smoother. Nevertheless, even with the implementation of our smoothing strategy, the CWT halo exhibits numerous sharp edges due to the limited resolution. Finally, we discard halos that contain too few particles and obtain our final identification result.
Figure 8: The largest halos in the four slices are identified by the CWT (top) and the FOF (bottom) methods. All particles within these halos are scattered and are colored according to their local density.

### v3

@@ 2.3 SIMBA Simulation and Pseudo-2D Data
SIMBA (Davé et al., 2019) is a meshless finite mass hydrodynamic cosmological simulation based on the GIZMO cosmological gravity plus hydrodynamics solver (Hopkins, 2015), assuming $h=0.68$, $\Omega_{\rm\Lambda}=0.7$, $\Omega_{\rm m}=\Omega_{\rm dm}+\Omega_{\rm b}=0.3$, $\Omega_{\rm b}=0.048$, $n_{s}=0.97$, and $\sigma_{8}=0.82$, which are the Planck cosmological parameters (Planck Collaboration et al., 2016). This simulation consists of several sub-simulations with different box lengths, particle numbers, and even different physics, e.g., no active galactic nuclei, no jet, and pure dark matter.
This paper aims to identify halos within the particle data, without delving into the detailed physics of the simulation. We therefore use the m50n512 dark matter sub-simulation sample, which is a pure dark matter simulation devoid of complicated processes such as star formation, gas cooling, and feedback [^2] This dataset was downloaded from http://simba.roe.ac.uk/simdata/m50n512/dm/snapshots/snap_m50n512_151.hdf5, but access to this sub-simulation is currently disabled.. This sub-simulation has a box length of $50h^{-1}\rm Mpc$ and contains $512^{3}$ dark matter particles, with a mass resolution of $9.6\times 10^{7}M_{\odot}$.
To obtain 2D data, we select the snapshot at redshift $z=0$ from this sub-simulation and cut a slice in the z-direction, keeping only the x, y coordinates of these particles. Though the dimension is reduced from the original three-dimensional (3D) data, this pseudo-2D dataset retains substantial structural information, proving adequate for testing our halo identification algorithm.
We use four slices of pseudo-2D data at ${\rm z}=11.4$, $28.4$, $38.0$ and $45.7\,{h^{-1}{\rm Mpc}}$ with a thickness of $\Delta{\rm z}=0.1\,{h^{-1}{\rm Mpc}}$, containing 323,716, 228,187, 303,692, and 312,372 particles, respectively. The CIC grids of these slices, each consisting of $512^{2}$ cells, are shown in Figure 1.
@@ 3.1 Procedure of Identification
(1) Calculation of the grid CWT at different scales using a grid-assigning method, as shown in Equation (6), which results in an $\mathcal{O}(N)$ time complexity.
(2) Determination of thresholds to filter those local maxima generated by Poisson noise at different scales, as discussed in Section 3.3, is performed using either the Monte Carlo method or a pre calculated empirical formula.
@@
In the following, we present the detailed calculations for each step of the process. A summary of the parameters utilized in this study is presented in Table 1.
Figure 2: The 2D CWT of the slice of ${\rm z}=38\,{h^{-1}{\rm Mpc}}$ with four different scale parameters, from top left to bottom right represent the CWT at $w=2.29$, $5.94$, $15.75$, and $35.24\,{h{\rm Mpc}^{-1}}$, respectively. The sizes of the structures presented in these CWTs vary depending on the scale parameter, extending from large to small scales.
Figure 3: The same halo is depicted with directly assigned particles (left) and weighted assigned particles (right). The color represents the probability density of particles. The halo with directly assigned particles exhibits a very serrated boundary, whereas the weighted assignment results in a smoother, more rounded boundary.
@@ 3.2 Grid-based CWT Computation
Equation (6) illustrates that the grid CWT is the sum of the wavelets centered on the position of the particles. Since the MH wavelet $\Psi(x)$ in Equation (8) has an approximate compact support, it affects only a finite range near its center. This means that for some $r_{\rm c}>0$, if $|\Psi(\mathbf{x})|$ is less than, e.g., $10^{-10}$ for all the $|\mathbf{x}|>r_{\rm c}$, then we can truncate $\Psi(|\mathbf{x}|)$ to be zero wherever $|\mathbf{x}|>r_{\rm c}$. This property allows a convenient calculation of the grid CWT, since a wavelet at any scale will only have a nonzero value at grid points near its center.
@@
In this way, $k_{w}=1$ always represents an intermediate scale when identifying halos, and in this study, it corresponds to a scale parameter of $20\,{h{\rm Mpc}^{-1}}$. A dynamic mesh is created based on the variable $k_{w}$ to save memory and speed up computation. To accurately compute these translated wavelets on the grid, the grid point closest to the position of the particle is selected as the center of the grid wavelet, and this grid point is extended by ${\rm Int}(10/k_{w})+10$ grid points in each direction. We will discuss the resolution parameters in the above two equations in detail in the next section.
The $N_{\rm g}^{2}$ grid is initialized with zeros, and each individual CWT, corresponding to its particle in the $N_{\rm W}^{2}$ grid, is calculated and added to the appropriate part of the $N_{\rm g}^{2}$ grid. This part is centered at the grid point nearest to the particle and extends ${\rm Int}(10/k_{w})+10$ grid points in each direction. If a particle is located near the edge of the simulation box, its extended ${\rm Int}(10/k_{w})+10$ grid points may overflow beyond the $N_{\rm g}^{2}$ grid. In this scenario, the overflowed grid points are relocated to the opposite side, using the periodic boundary condition applied to the particle data.
@@ 3.3 Cross-scale Analysis of Local Maxima and Thresholding
Once the CWT grid is computed, the clustering of particles can be easily analyzed (Slezak et al., 1990). When using wavelets to identify haloes, it is generally agreed that each halo should be located at the position of its corresponding local maximum. The scale at which we find the maximum represents the size of the structure (e.g., Bendjoya et al., 1991; Flin & Krywult, 2006; Grebenev et al., 1995; Slezak et al., 1993; Cayón et al., 2000; Lazzati et al., 1999; Patrikeev et al., 2006; Hayn et al., 2012). Since our wavelet is not orthogonal, finding its local maximum directly in 3D space (2D location plus 1D-scale parameter) can be misleading, as a single structure may reach its maximum at multiple scales (Hayn et al., 2012; Ellien et al., 2021). Consequently, it is imperative to verify these maxima across scales to ensure that we identify the maximum that corresponds to the actual structure, that is, the most prominent maximum near the position of this structure.
@@
Also, the particle distribution represents a Poisson sampling of the so-called real density field. This sampling process introduces Poisson noise that can lead to spurious local maxima when calculating the CWT using these particle data (Bijaoui et al., 1992). To mitigate this issue, a significance test should be conducted to assess the likelihood of detecting true structures as opposed to Poisson noise. We apply thresholds obtained from Monte Carlo simulations and eliminate any maxima that do not meet these thresholds across each scale. A detailed description of this process will be presented in the subsequent section.
@@
Once we have identified all thresholded local maxima, we divide the simulation box into an additional $N_{\rm cs}^{2}$ grid. $N_{\rm cs}$ represents the grid resolution for the cross-scale comparison and we set its value to 1000 as summarized in Table 1. Due to the use of the dynamic mesh in the calculation of the grid CWT, a consistent spatial resolution is required to locate these maxima when comparing them across scales.
@@
We then convert the location of these maxima, which is expressed as the count of grid lengths at their corresponding scale, onto the $N_{\rm cs}^{2}$ grid, starting from the larger to the smaller physical scales (from a smaller $w$ to a larger $w$). When we map a maximum onto the $N_{\rm cs}^{2}$ grid for cross-scale comparison, we compare its CWT value with the values of the surrounding $3\times 3$ grid points centered on the grid point of this maximum. If this maximum has a higher CWT value than all nine values within the surrounding grid points, we clear the information stored in those grid points and record the information of this maximum in its corresponding grid point. This ensures that there is only one maximum with the highest CWT value within every $3\times 3$ grid block of the $N_{\rm cs}^{2}$ grid. These maxima correspond to halo candidates in our method.
@@ 3.4 Grid Segmentation and Halo Boundary Delineation
To identify a halo, a clear boundary must be established between the halo and the background Universe. Fortunately, the wavelet method provides a natural boundary through its inherent properties. A positive CWT value is obtained in regions where the signal’s shape aligns with that of the wavelet, such as at the signal’s peak. Conversely, a negative CWT value is observed in regions where the signal’s shape is inversely related to the wavelet, such as within the signal’s valley. Additionally, the matching between the peak and the negative ring of our wavelet also leads to negative CWT values. In Figure 2 we present four different CWTs for the slice at ${\rm z}=38.0\,{h^{-1}{\rm Mpc}}$, covering physical scales from large to small. The result confirms the ability of the CWT to distinguish structures across scales. The CWT at a larger scale contains larger clumps, and as the scale parameter increases (toward a smaller physical scale), the size of these clumps decreases. The CWT at any given scale reflects the clustering of particles into clumps at a size corresponding to this scale, and the value of the CWT represents the amplitude or strength of its corresponding clump.
Notably, all four CWTs exhibit a spatial distribution similar to that of the CIC density shown in Figure 1. Higher densities in the density field also result in higher CWT values, while the absence of particles corresponds to a CWT value approaching zero. Additionally, every positive clump in the CWT is encircled by a high-negative ring caused by the matching of the negative ring of the wavelet and the density peak. If a wavelet matches with a peak in a positive signal, it indicates that the positive part has a size similar to that of the peak, with the negative ring located outside. The location where a positive CWT value is attained is situated within the density peak, whereas the exterior portion, where the CWT value is negative, is situated outside of the density peak. In this way, points where the CWT equals zero are intermediate positions and serve as the boundary of the peak in the CWT field. However, since the CWT is calculated on a finite grid, finding the grid points where the CWT equals zero can be insufficient. In some cases, there may not be any grid points that are strictly equal to zero.
@@
Fortunately, we can also use all the outer regions as boundaries to isolate halos. This allows us to segment the full CWT grid at each scale into ‘islands’ of positive values surrounded by a ‘sea’ of zero or negative values. These islands represent clusters at that scale. Similar segmentation methods have been employed in many previous studies to identify various structures within the CWT grid (e.g., Baluev & Rodionov, 2020; Freeman et al., 2002), as well as in the DWT grid (e.g., Pagliaro et al., 1999; Mertens & Lobanov, 2015).
@@
This method proves effective for small and medium physical scales but encounters limitations at larger physical scales. The support of our wavelet is more extensive at larger scales, which results in the linking of nearby structures by ‘positive bridges’. Additionally, the lower grid resolution at larger scales exacerbates this situation. Regardless of the constant boundary condition applied, some structures are invariably linked by a positive bridge.
@@
Despite the presence of substructures, a gravity-bound halo exhibits a unimodal structure. At a larger scale, these substructures have a negligible impact on the CWT. Therefore, any actual halo must have a unimodal CWT. With this in mind, we introduce a new boundary condition akin to that described in Freeman et al. (2002), which divides closely situated structures by identifying the saddle point between them. At any scale, we start by identifying all local maxima in the CWT and use these as the basis for initial segmentation. These maxima grid points are ‘seeds’ of structures, and we expand them in an iterative process to obtain the final structures. During each iteration, we assess all connected grid points to ascertain whether any are lower than their connected structure grid points but higher than zero. We then expand these structures to include neighboring grid points that meet the aforementioned criteria. The expansion of structures continues until no neighboring grid points can be incorporated. At that point, our segmentation process is complete.
@@
As mentioned in Section 3.3, a single structure may manifest as multiple islands across different scales. Therefore, it is necessary to associate each halo candidate maximum with an island to avoid duplicate detections. Furthermore, it is crucial to exclude substructures, as they might also manifest as local maxima and islands, which can lead to confusion when identifying halos. We zoom in on all the islands and their corresponding background sea at different scale parameters to the highest resolution, which is the resolution of the largest scale parameter $w$. Every halo candidate is then assessed from smaller to larger scale parameters. If a halo candidate is located within an island at its scale and this island does not overlap with any confirmed structures, it is registered as a halo.
Figure 4: Left: threshold-$k_{w}$ relationship at different particle numbers, ranging from $N=10^{5}$ (bottom, black) to $10^{7}$ (top, yellow). We can observe that the relationship between the variables demonstrates high linearity. Right: the slope and intercept of the linear relation from the left panel are shown as functions of the particle number. The dotted lines represent the empirical formulas fitted to the data.
@@ 3.5 Refinement and Smoothing of Halo Boundaries
After registering all halos on a map with the highest resolution (smallest scale), we assess each particle to determine if it is located within a halo and label it with the corresponding halo index. However, constrained by our grid resolution, the CWT halo in the densest regions has very sharp edges. To address this issue, we propose checking if the particle is located within any particular halo grid, as shown in the left panel of Figure 3, where the color encodes the probability density of particles derived by a Gaussian kernel. To alleviate the serrated edges, we examine the halo attribution of the four nearest grid points for each particle and calculate the particle’s weight to the four grid points using a CIC weighting formula, as
@@
where ${\rm x}_{i}$, ${\rm y}_{j}$ represent the normalized distances to the $i,j$ th grid point in x and y direction, respectively, and $i,j\in[1,2]$ denote the four nearest grid points. After obtaining the weights, we sum them up if some grid points are assigned to the same halo. The particle is then assigned to the halo with the largest weight (we consider the background as halo ‘0’). As illustrated in Figure 3, the once severe serrated edges have become noticeably smoother following the weighting process. Despite our smoothing strategy, the CWT halos still present numerous sharp edges, a consequence of the limited resolution. In the 3D case, these sharp edges can be further suppressed through an unbinding procedure as is the case with most halo finders, because we can access the full dynamical information of those particles instead of limited pseudo-2D data. Finally, we discard halos that contain too few particles and obtain our final identification result.